# TimeRecipe: A Time-Series Forecasting Recipe via Benchmarking Module Level Effectiveness

### v1

@@ B.1 Additional Dataset Explanation
ETT, alias Electricity Transformer Temperature [86], contains collected oil temperature and electricity load data per minute of two Chinese stations between 2016/07 to 2018/07. The original data is then aggregated every one hour (ETTh1/2) or every 15 minutes (ETTm1/2).
@@
ECL, alias Electricity Consumption Load [7], contains the hourly electricity consumption history of 321 clients covering two years.
@@
Exchange [31] contains daily exchange rates of U.S. dollars in eight foreign countries, including Australia, Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore, between 1990 and 2016.
@@
Our collected data covers multiple domains, multiple granularity, and multiple time ranges, ensuring plenty of diversity like the corresponding series. By evaluating such datasets, we guarantee that the outputted recipe from our method maintains considerable coverage in real-world time-series tasks.
@@ B.2 Evaluation Metric Formulation
We follow the common evaluation metrics across diverse forecasting datasets [55, 39, 69]. We evaluate the LTSF dataset and PEMS dataset by two metrics: mean squared error (MSE) and mean absolute error (MAE), aiming to evaluate the errors under the same scale. Given a predicted sequence $\hat{\mathbf{Y}}=\{\hat{x}_{t+1},\dots,\hat{x}_{t+H}\}$ and corresponding ground truth $\mathbf{Y}=\{x_{t+1},\dots,x_{t+H}\}$ where $\mathbf{Y},\hat{\mathbf{Y}}\in\mathbb{R}^{H\times d_{X}}$, these two metrics are calculated by:
@@ B.3 Dataset Properties Measurement
We define the data properties involved in TimeRecipe following TFB [56].
Trend refers to the long-term changes of time-series. As shown in Algorithm 1, we involve Seasonal and Trend decomposition using Loess (STL) [14] as a decomposition function to calculate the trending values.
@@
Correlation refers to the possibility that different channels of a multivariate sequence share a similar distribution. As shown in algorithm 6, we calculate the correlation referring to Catch22 [47] library, which is designed to extract 22 features from time series.
@@ Appendix C Detailed Results
C.1 Comprehensive Claims (Section 4.1.2 Cont.)
Here we present all the claims on the correlations between the effectiveness of the module and the data characteristics that are statistically significant (e.g., $\leq$0.05). The comprehensive results are shown in Table 3.
@@
C.2 Performance Ranking Details (Section 4.1.1 Cont.)
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type SMAPE Rank MASE Rank OWA Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
@@ Appendix D Limitation Discussion
Not all time-series designs are covered. We recognize that TimeRecipe does not include every existing module design in time-series forecasting. The designs not covered generally fall into two categories. First, there are highly specialized designs, such as those in Crossformer [82]. These are specifically created for the Transformer architecture to capture dependencies across time and dimensions but are difficult to apply to other model architectures. Second, some model architectures are not yet widely adopted, such as models based on large language models (LLMs), for example, Time-LLM [26]. Such models often have high computational costs, and their effectiveness is still a subject of ongoing discussion [63].

### v2

@@ B.1 Additional Dataset Explanation
ETT, alias Electricity Transformer Temperature (Zhou et al., 2021), contains collected oil temperature and electricity load data per minute of two Chinese stations between 2016/07 to 2018/07. The original data is then aggregated every one hour (ETTh1/2) or every 15 minutes (ETTm1/2).
@@
ECL, alias Electricity Consumption Load (Asuncion et al., 2007), contains the hourly electricity consumption history of 321 clients covering two years.
@@
Exchange (Lai et al., 2018) contains daily exchange rates of U.S. dollars in eight foreign countries, including Australia, Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore, between 1990 and 2016.
@@
Our collected data covers multiple domains, multiple granularities, and multiple time ranges, ensuring plenty of diversity like the corresponding series. By evaluating such datasets, we guarantee that the outputted recipe from our method maintains considerable coverage in real-world time-series tasks.
@@ B.2 Evaluation Metric Formulation
We follow the common evaluation metrics across diverse forecasting datasets (Oreshkin et al., 2019; Liu et al., 2022a; Wang et al., 2024a). We evaluate the LTSF dataset and PEMS dataset by two metrics: mean squared error (MSE) and mean absolute error (MAE), aiming to evaluate the errors under the same scale. Given a predicted sequence $\hat{\mathbf{Y}}=\{\hat{x}_{t+1},\dots,\hat{x}_{t+H}\}$ and corresponding ground truth $\mathbf{Y}=\{x_{t+1},\dots,x_{t+H}\}$ where $\mathbf{Y},\hat{\mathbf{Y}}\in\mathbb{R}^{H\times d_{X}}$, these two metrics are calculated by:
@@ B.3 Dataset Properties Measurement
We define the data properties involved in TimeRecipe following TFB (Qiu et al., 2024).
Trend refers to the long-term changes of time-series. As shown in Algorithm 1, we involve Seasonal and Trend decomposition using Loess (STL) (Cleveland et al., 1990) as a decomposition function to calculate the trending values.
@@
Correlation refers to the possibility that different channels of a multivariate sequence share a similar distribution. As shown in algorithm 6, we calculate the correlation referring to Catch22 (Lubba et al., 2019) library, which is designed to extract 22 features from time series.
@@ Appendix C Detailed Results
C.1 Comprehensive Claims (Section 4.1.2 Contd.)
[…]
@@
C.2 Model Selection Toolkit Evaluation (Section 4.2 Contd.)
Here, we present additional evaluation results for our training-free model selection approach on ETTh1_24_M, Environment_48_S, and Security_12_S. As shown in Table 5, these results are consistent with the conclusions drawn from Table 3.
ETTh1_24_M
IN SD Fusion Embed $F$-Type Rank MSE MAE
TimeRecipe Best Feature Patch MLP 1.5 0.2963 - 0.3467 -
Top-3 Selection Feature Patch MLP 1.5 0.2963 0.0% 0.3467 0.0%
Temporal Patch MLP 20.5 0.3096 -4.3% 0.3586 -3.4%
Temporal Patch Trans. 4.0 0.2988 -0.9% 0.3520 -1.5%
Selection better than existing best: Yes (Selection 1&3)
Environment_48_S
IN SD Fusion Embed $F$-Type Rank MSE MAE
TimeRecipe Best Temporal Patch RNN 2.0 0.2912 - 0.3786 -
Existing Best (DLinear) Temporal None MLP 16.0 0.2950 -1.3% 0.3825 -1.0%
Top-3 Selection Feature Patch RNN 21.5 0.2987 -2.5% 0.3824 -1.0%
Feature Patch RNN 36.5 0.3068 -5.3% 0.3936 -4.0%
Feature Token RNN 2.5 0.2920 -0.2% 0.3781 +0.1%
Selection better than existing best: Yes (Selection 3)
Security_12_S
IN SD Fusion Embed $F$-Type Rank MSE MAE
TimeRecipe Best Feature None RNN 2.5 74.2170 - 4.0465 -
Existing Best (DLinear) Temporal None MLP 26.0 75.2914 -1.4% 4.3136 -6.6%
Top-3 Selection Feature Invert Trans. 56.0 86.9113 -17.1% 5.1710 -27.8%
Temporal Patch RNN 19.0 75.2749 -1.4% 4.2357 -4.7%
Temporal Patch MLP 14.5 74.9425 -0.9% 4.2124 -0.4%
Selection better than existing best: Yes (Selection 2&3)
Table 5: Comparison of model configurations selected by our method against the best-performing existing model and the global best results found through exhaustive design space search. At least one of our top-3 selections consistently outperforms the existing best and closely approaches the optimal configuration on ETTh1_24_M, Environment_48_S, and Security_12_S. The results demonstrate the value of TimeRecipe, even when using a simple tree-based approach.
C.3 Performance Ranking Details (Section 4.1.1 Contd.)
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type SMAPE Rank MASE Rank OWA Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@
Setup IN SD Fusion Embed $F$-Type MSE Rank MAE Rank Total
@@ Appendix D Limitation Discussion
Not all time-series designs are covered. We recognize that TimeRecipe does not include every existing module design in time-series forecasting. The designs not covered generally fall into two categories. First, there are highly specialized designs, such as those in Crossformer (Zhang & Yan, 2023). These are specifically created for the Transformer architecture to capture dependencies across time and dimensions but are difficult to apply to other model architectures. Second, some model architectures are not yet widely adopted, such as models based on large language models (LLMs), for example, Time-LLM (Jin et al., ). Such models often have high computational costs, and their effectiveness is still a subject of ongoing discussion (Tan et al., 2024).